below o A double coer utePI(int eurber) 1) return
Answer #1

The answer is B. Given a number 'n' we need to calculate the sum of the following series

\frac{1}{1^{2}} + \frac{1}{2^{2}} +\frac{1}{3^{2}} +... +\frac{1}{n^{2}}
which can further be simplified into the following form
\frac{1}{1*1} + \frac{1}{2*2} +\frac{1}{3*3} +... +\frac{1}{n*n}

To get the sum of the above series, we need to first find compute the value of \frac{1}{n*n} and add it to the sum of the \frac{1}{1^{2}} + \frac{1}{2^{2}} +\frac{1}{3^{2}} +... +\frac{1}{(n-1) * (n-1)} , which is upto (n-1)th term.

Now, to get the sum of the above series, we need to first compute the value of \frac{1}{(n-1) * (n-1)} and it to the sum of the remaining series \frac{1}{1^{2}} + \frac{1}{2^{2}} +\frac{1}{3^{2}} +... +\frac{1}{(n-2) * (n-2)} , which is upto (n-2)th term. And so on, we compute the value of the entire series, in a recursive manner.

The following function does the same thing as discussed above. It first computes the value of 1/(n * n) and add it to the sum of the series upto (n-1)th term, computed in a recursive manner. Each time you go on decrementing n by 1 until you reach 1. At this point you will return 1 and all the terms starts up getting added.

double computePI(int number) {

if (number <= 1) {
          return 1.0; // This is the base case
}
return 1 / (number * number ) + computePI(number - 1);

}